### v1

# Elliptic functions, Floquet transform and Bergman spaces on doubly periodic domains

Abstract.
We study Bergman spaces $A^{2}(\Omega)$, their kernels and Toeplitz operators on unbounded, doubly periodic domains $\Omega$ in the complex plane. We establish the mapping properties of the Floquet transform operator defined in $A^{2}(\Omega)$ and derive a general formula connecting the Bergman kernel and projection of the domain $\Omega$ to a kernel and projection on the bounded periodic cell $\varpi$. As an application, we prove, for Toeplitz operators $T_{a}$ with doubly periodic symbols, a spectral band formula, which describes the spectrum and essential spectrum of $T_{a}$ in terms of the spectra of a family of Toeplitz-type operators on the cell $\varpi$.
Technical challenges arise from the fact that double quasiperiodic boundary conditions have to be taken into account in the definitions of the spaces and operators on the periodic cell $\varpi$. This requires novel operator theoretic tools, which are based on modifications of certain elliptic functions, e.g. the Weierstrass $\wp$-function.
The research of the first named author was partially supported by the Academy of Finland grant no. 359563
1. Introduction.
In this paper we consider Bergman space $A^{2}(\Omega)$ and the Bergman projection $P_{\Omega}$ on planar domains $\Omega\subsetneq{\mathbb{C}}$, which have the special geometry of being periodic in two directions. For simplicity, we require that
(1.1) $z\in\Omega\ \Rightarrow\ \ z+m\in\Omega\ \mbox{for all}\ m\in\Lambda=\{z\in{\mathbb{C}}\,:\,{\rm Re}\,z\in{\mathbb{Z}},\ {\rm Im}\,z\in{\mathbb{Z}}\}.$
Following the paper [10], where the singly periodic case was treated, the aim is to apply Floquet transform techniques to describe the connection of the Bergman projection and kernel on $\Omega$ with projections and kernels on the periodic cell $\varpi$. We will apply these considerations to study the spectra of Toeplitz operators with doubly periodic symbols.
Given a domain $\Omega$ in the complex plane $\mathbb{C}$}, we denote by $L^{2}(\Omega)$ the usual Lebesgue-Hilbert space with respect to the (real) area measure dAdA and by $A^{2}(\Omega)$ the corresponding Bergman space, which is the closed subspace consisting of analytic functions. The Floquet transform is defined for $f\in A^{2}(\Omega)$ by
(1.2) ${\sf F}f(z,\eta)$ $=$ $\widehat{f}(z,\eta)=\frac{1}{2\pi}\sum_{m\in\Lambda}e^{-i\eta_{1}m_{1}-i\eta_{2}m_{2}}f(z+m),$
where $z\in\varpi$ and $\eta=(\eta_{1},\eta_{2})\in{\bf Q}=[-\pi,\pi]^{2}$ is the so called Floquet-parameter or quasimomentum; here and later, it will be convenient to denote $m_{1}={\rm Re}\,m$, $m_{2}=\ {\rm Im}\,m$ for a number $m\in\Lambda$. By using the canonical identification of ${\mathbb{C}}\cong{\mathbb{R}}^{2}$, this definition coincides with the standard definition of the Floquet transform in periodic subdomains of ${\mathbb{R}}^{2}$. In particular, the well-known properties of Fourier series (see also [6], Section 4.2) imply that $\sf F$} is a continuous (isometric) mapping from $L^{2}(\Omega)$ onto $L^{2}({\bf Q};L^{2}(\varpi))$, which is the vector valued $L^{2}$ space, or the space of $L^{2}$-Bochner integrable functions $g:{\bf Q}\to L^{2}(\varpi)$; see [3]. If $g\in L^{2}({\bf Q};L^{2}(\varpi))$, we also denote
(1.3) ${\sf F}^{-1}g(z)$ $=$ $\frac{1}{2\pi}\int\limits_{{\bf Q}}e^{i[{\rm Re}z]\eta_{1}+i[{\rm Im}z]\eta_{2}}g(z-[{\rm Re}\,z]-i[{\rm Im}z],\eta)d\eta,\ \ z\in\Omega.$
We refer to [5], [6] for an introduction of the Floquet transform in the study of the Schrödinger equation with periodic potentials and to [7], [8] for its use in periodic elliptic spectral problems. See also [10] for a slightly more thorough introduction to the topic in the setting of analytic function spaces. Taking the restriction of $\sf F$} to the subspace of analytic functions, we conclude that $\sf F$} is a unitary mapping from $A^{2}(\Omega)$ into $L^{2}({\bf Q};L^{2}(\varpi))$, but in order to achieve the bijectivity, there remains to determine the image ${\sf F}(A^{2}(\Omega))$ in the Bochner space. In [10], the first named author established the mapping properties of $\sf F$} and solved this problem in the case of complex domains $\Pi$, which are periodic, say, in the direction of the real axis and bounded in the imaginary direction and thus contained in a strip.
The aim of this paper is to extend many of the results of [10], [11] to the doubly periodic case, but having a closer look at the arguments of these papers, one observes that two essential technical tools of the citations fail in the present setting. To characterize the range of the Floquet transform in the Bergman space case on singly periodic domains, [10], it was necessary to approximate a Bergman function $f\in A^{2}(\Pi)$ by a function $\varphi_{\varepsilon}f$, where $\varphi_{\varepsilon}(z)=e^{-\varepsilon z^{2}}$, $z\in\Pi$, $\varepsilon>0$, is a rapidly decreasing function as $|z|\to\infty$ in $\Pi$. Indeed, the function $\varphi_{\varepsilon}$ has Gaussian decay at infinity, if $z$ belongs to a strip parallel to the real axis, but loses this property, if $z$ is allowed to belong to a doubly periodic domain, which is automatically unbounded in the imaginary direction. The same problem appears when trying to generalize the results of [11], where the operator $J_{\nu,\mu}:f\mapsto e^{i(\nu-\mu)z}f$ was used to switch the Floquet parameters $\nu,\mu\in[-\pi,\pi]$ in the quasiperiodic boundary conditions.
In Section 2 we construct, by using the theory of elliptic (doubly periodic meromorphic) functions, the tools that allow us to adapt the methods of [10], [11] to the doubly periodic case. These consist of the functions $\psi^{(\rho)}$ and $\psi_{\eta}$, see Lemmas 2.4 and 2.6, which replace the functions $e^{-\varepsilon z^{2}}$ and $e^{i\eta z}$ used in the singly periodic case.
The remaining sections are devoted to the formulation of the main results and their proofs. In Section 3, Theorem 3.4, we characterize the image ${\sf F}(A^{2}(\Omega))$ in $L^{2}({\bf Q};L^{2}(\varpi))$ and in Theorem 4.1 of Section 4 we present the relation of the Bergman kernels in the domains $\Omega$ and $\varpi$. In Section 5 we apply the results of the preceding sections to study spectra of Toeplitz operators $T_{a}$ with periodic symbols $a$ on the space $A^{2}(\Omega)$. Theorem 5.1 contains a proof for the spectral band formula, which characterizes the spectrum and essential spectrum of $T_{a}$ in terms of a family of spectra of $\eta$-dependent Toeplitz-type operators in the periodic cell.
We denote by $P_{\Omega}$ the orthogonal projection from $L^{2}(\Omega)$ onto $A^{2}(\Omega)$. It can always written with the help of the Bergman kernel $K_{\Omega}:\Omega\times\Omega\to{\mathbb{C}}$,
(1.4) $P_{\Omega}f(z)=\int\limits_{\Omega}K_{\Omega}(z,w)f(w)dA(w)$
and the kernel has the properties that $K_{\Omega}(z,\cdot)\in L^{2}(\Omega)$ for all $z$ and $K(z,w)=\overline{K(w,z)}$ for all $z,w\in\Omega$. See e.g. [4] for a proof of these assertions.
The basic assumptions on the domain $\Omega$ and its periodic cells are as follows. However, we will pose one more assumption, related to the theory of elliptic functions, see $({\mathcal{A}})$ in the next section. Here and later, cl(A)(A) denotes the closure of the set AA.
Definition $1.$
We denote by $\varpi$ the periodic cell, which is a domain in $\mathbb{C}$} such that
$(i)$ $\varpi\subset Q=(0,1)\times(0,1)\subset{\mathbb{R}}^{2}\cong{\mathbb{C}},$
$(ii)$ the boundary $\partial\varpi$ contains the boundary $\partial Q$, and
$(iii)$ the complement $Q\setminus{\rm cl}(\varpi)$ contains an open set
We denote the translates of $\varpi$ by $\varpi_{m}=\varpi+m$ for all $m\in\Lambda$. The periodic domain $\Omega$ is defined as the interior of the set
$\bigcup_{m\in\Lambda}{\rm cl}(\varpi_{m}).$
We emphasize that contrary to [10], we do not need to add assumptions on the smoothness of the boundary $\partial\Omega$. Anyway, the periodic domain $\Omega$ is always infinitely connected, due to assumption $(iii)$ of Definition $1.$
The following notation will be used throughout the paper. We write $C$, $C^{\prime}$, $\ldots$, (respectively, $C_{\eta}$, $C^{\prime}_{\eta}$, $\ldots$ etc.) for positive constants independent of the functions or variables in the given inequalities (resp. depending only on a parameter $\eta$ etc.), the values of which may vary from place to place. For $\in{\mathbb{R}}, [x][x$] denotes the largest integer not larger than $x$. We write ${\mathbb{Z}}=\{0,\pm 1,\pm 2,\ldots\}$. Subsets of the complex plane will often be described using the real planar coordinates, for example, rectangles in $\mathbb{C}$} are described as the sets $a,b]\times[c,d]$. If $x$ is a point in $\mathbb{C}$} or ${\mathbb{R}}^{n}$ and $r>0$, then $B(x,r)$ denotes the Euclidean ball with center $x$ and radius $r>0$. Moreover, cl(A)(A) denotes the closure of a set AA.
Given a domain $D\subset{\mathbb{C}}$, we denote by $\|f\|_{D}$ and $(\cdot|\cdot)_{D}$ the norm and inner product of $L^{2}(D)$, which is the $L^{2}$ space with respect to the (real) area measure. In general, the norm of a Banach space $X$ is denoted by $\|\cdot\|_{X}$. Given an interval $I\subset{\mathbb{R}}$ and a Banach space $X$ we denote by $L^{2}(I;X)$ the space of vector valued, Bochner-$L^{2}$-integrable functions on II with values in $X$, endowed with the norm
$\|f\|_{L^{2}(I;X)}=\Big(\int\limits_{I}\|f(t)\|_{X}dt\Big)^{1/2}\,.$
If $X=L^{2}(D)$ for some domain $D$, then $L^{2}(I;X)$ is a Hilbert space endowed with the inner product $\int_{I}(f(t)|g(t))_{D}dt$. See [3] for the theory of Bochner spaces.
Given a Banach space $X$, ${\mathcal{L}}(X)$ stands for the Banach space of bounded linear operators $X\to X$. The operator norm of $T\in{\mathcal{L}}(X)$ is denoted just by $\|T\|$ or by $\|T\|_{X\to X}$, if it is necessary to specify the domain or target spaces. The identity operator on $X$ is written as $I_{X}$ or by $I_{D}$, if $X=L^{2}(D)$. For an operator $T\in{\mathcal{L}}(H)$, where $H$ is a Hilbert space, $\sigma(T)$, $\sigma_{\rm ess}(T)$ and $\varrho(T)$ stand for the spectrum, essential spectrum and resolvent set of the operator $T$. The resolvent (operator) of $T\in{\mathcal{L}}(H)$ is denoted by $R_{\lambda}(T)=(T-\lambda I_{H})^{-1}$, where $\lambda\in\varrho(T)$.
2. On elliptic functions in the periodic domain $\Omega$.
Our aim in this section is to prove Lemmas 2.4 and 2.6 which provide the technical tools needed for the proofs of the main results in later sections. The tools consist of modifications of certain elliptic functions. Recall that an elliptic function $\varphi$ is by definition a meromorphic and doubly periodic function in the plane $\mathbb{C}$}, see for example [2], Chapter 5, or [9], Chapter 14, and others. It ought to be recalled that, except for the constant functions, there does not exist doubly periodic functions which are analytic in the entire complex plane (Theorem 1, Chapter 14 of [9]), and the use of meromorphic ones is thus necessary for our purposes. Here, in view of our choice of the domain $\Omega$, we will assume the periodicity
(2.1) $\varphi(z)=\varphi(z+1)=\varphi(z+i).$
According to the classical theory, any elliptic function satisfying (2.1) has a finite number of zeros in the cell $0,1)\times[0,1)\subset{\mathbb{C}}$ and equally many poles there as well, when counted by taking into account the zero and pole orders (Theorem 5, Chapter 14 of [9])
Let us make the following assumption on the domain $\Omega$.
$({\mathcal{A}})$ There exists an elliptic function $\varphi$ with periodicity (2.1), such that the zeros of $\varphi$ occur at points $z\in{\mathbb{C}}\setminus{\rm cl}(\Omega)$ and such that $\varphi$ has a representation
(2.2) $\varphi(z)=\sum_{m\in\Lambda}\phi(z-m)$
for some function $\phi$, which is meromorphic in $\mathbb{C}$} and has only finitely many poles, all situated in $Q$ and which satisfies
(2.3) $\sup_{z\in\Omega\setminus{\rm cl}(Q)}(1+|z|)^{b}|\phi(z)|=:d_{\phi}<\infty$
for some constant $b>2$.
We denote by $S\subset Q$ the finite set of the poles of the function $\phi$. Note that (2.3) implies the absolute convergence of the series (2.2), and there actually holds a stronger statement, see Lemma 2.3.
Example 2.1.
We show that $({\mathcal{A}})$ holds, if $Q\setminus{\rm cl}(\varpi)$ contains the three points $\alpha+\frac{1}{2}$, $\alpha+\frac{i}{2}$ and $\alpha+\frac{1+i}{2}$ for some $\alpha$ with Re $\alpha\in(0,\frac{1}{2})$ and Im $\alpha\in(0,\frac{1}{2})$. Indeed, the Weierstrass $\wp$-function is defined by
(2.4) $\wp(z)=\frac{1}{z^{2}}+\sum_{m\in\Lambda\setminus\{0\}}\Big(\frac{1}{(z-m)^{2}}-\frac{1}{m^{2}}\Big),$
and its well-known properties include the following ([9], Chapter 14) :
$(i)$ $\wp$ is a meromorphic function in $\mathbb{C}$} with poles of order 2 at the points $m\in\Lambda$,
$(ii)$ the zeros of $\wp$ occur at the points $\frac{1}{2}(1+i)+m$, $m\in\Lambda$, and
$(iii)$ $\wp$ is doubly periodic with periodicity (2.1).
The representation (2.4) is not of the form (2.2), but the derivative $\wp^{\prime}$ can be calculated by termwise differentiation, which yields
$\wp^{\prime}(z)=\sum_{m\in\Lambda}\frac{-2}{(z-m)^{3}},$
i.e., a function of the form (2.2). It is known that the zeros of $\wp^{\prime}$ occur at the points $\frac{1}{2}$, $\frac{i}{2}$ and $\frac{1}{2}(1+i)$ and their translates, see the references above, or [1], Section 1. We define $\phi(z)=-2(z-\alpha)^{-3}$ and $\varphi(z)=\wp^{\prime}(z-\alpha)$, where $\alpha$ is a above. It is plain that with these functions, the domain $\Omega$ satisfies condition $({\mathcal{A}})$.
From now on we assume that $\Omega$ is a doubly periodic domain satisfying condition $({\mathcal{A}})$ with functions $\varphi$ and $\phi$. In view of Example 2.1, this is only a minor restriction of generality.
Remark $2.$
The function $\varphi$ is bounded from below on $\Omega$: there exists a constant $C>0$ such that $\inf_{z\in\Omega}|\varphi(z)|\geq C$. Namely, by the basic assumptions, the zeros of $\varphi$ are outside the set ${\rm cl}(\Omega)$. The claim follows from the periodicity of $\varphi$, which allows us to restrict the proof of the claim into a compact subset of $\mathbb{C}$}.
We will need two modifications of the function $\varphi$. The first one is given in Lemma 2.4 and it gives a replacement $\psi^{(\rho)}$ of the function $\varphi_{\varepsilon}$ of the singly periodic case (see Section 1). We first consider the following statement.
Lemma 2.3.
If $K$ is a compact set contained in ${\mathbb{C}}\setminus\bigcup_{m\in\Lambda}(S+m)$, then, for every $\delta\in(0,b-2)$,
(2.5) $\sum_{m\in\Lambda}|m|^{b-2-\delta}\sup_{z\in K\cap\Omega}|\phi(z-m)|<\infty.$
Proof. Given a compact $K$ as in the assumption, there are constants $C,C^{\prime}>0$ such that $|m|\leq C|z-m|$ for all $z\in K$, if $|m|\geq C^{\prime}$. We obtain
(2.6) $\sum_{\stackrel{{\scriptstyle\scriptstyle m\in\Lambda}}{{|m|\geq C^{\prime}}}}|m|^{b-2-\delta}\sup_{z\in K\cap\Omega}|\phi(z-m)|=\sum_{\stackrel{{\scriptstyle\scriptstyle m\in\Lambda}}{{|m|\geq C^{\prime}}}}|m|^{-2-\delta}\sup_{z\in K\cap\Omega}|m|^{b}|\phi(z-m)|$
$\leq$ $C\sum_{\stackrel{{\scriptstyle\scriptstyle m\in\Lambda}}{{|m|\geq C^{\prime}}}}|m|^{-2-\delta}\sup_{z\in K\cap\Omega}|z-m|^{b}|\phi(z-m)|$
By (2.3), we have here
(2.7) $|z-m|^{b}|\phi(z-m)|\leq d_{\phi}$
for all $z\in K$ except possibly for $z$ belonging to the cell $\varpi_{m}$, since $\phi$ has poles in $Q$. However, if $z\in\varpi_{m}\cap K$, then $|z-m|\leq 1$ and $|\phi(z-m)|$ is still bounded by a constant depending on $K$ only, since we are assuming that the distance of $K$ from the set $S+m$ is positive. Hence, (2.7) holds for all $K$, with possibly a larger constant on the right hand side. We conclude that (2.6) is bounded by
$C\sum_{m\in\Lambda}|m|^{-2-\delta}<\infty.$
The remaining finitely many terms of (2.5) with $|m|<C^{\prime}$ are uniformly bounded, by the assumption of the lemma. $\Box$
The construction and important properties of $\psi^{(\rho)}$ are contained in the next lemma. It uses the elliptic function $\varphi$ associated with the domain $\Omega$. Recall that $b>2$ was determined in (2.3).
Lemma 2.4.
Let $\varphi$, $\phi$ be as in (2.2), let $\rho\in(0,1]$ and define the functions
(2.8) $\varphi^{(\rho)}(z)=\sum_{m\in\Lambda}2^{-\rho|m|}\phi(z-m),\ z\in\Omega,\ \ \ \mbox{and}\ \ \ \psi^{(\rho)}=\frac{\varphi^{(\rho)}}{\varphi}.$
$(i)$ For every $\rho\in(0,1]$, the functions $\varphi^{(\rho)}$ and $\psi^{(\rho)}$ are meromorphic in $\mathbb{C}$}, and $\psi^{(\rho)}$ is analytic in $\Omega$.
$(ii)$ There exists a constant $C>0$ such that
(2.9) $\sup_{z\in\Omega,\rho\in(0,1]}|\psi^{(\rho)}(z)|\leq C.$
Moreover, for every $M\geq 1$, there holds $\psi^{(\rho)}\to 1$ uniformly on $\big([-M,M]\times[-M,M]\big)\cap\Omega$ as $\rho\to 0$.
$(iii)$ Given $f\in L^{2}(\Omega)$, there also holds
(2.10) $\|f\psi^{(\rho)}-f\|_{\Omega}\to 0\ \ \ \mbox{as}\ \rho\to 0.$
$(iv)$ For all compact $K\subset{\mathbb{C}}$ and $m\in\Lambda$, $\rho\in(0,1]$, we have
(2.11) $\sup_{z\in K\cap\Omega}|\psi^{(\rho)}(z+m)|\leq\frac{C_{K}}{\rho^{b}(1+|m|)^{b}}.$
Proof. $(i)$ Let us fix $\rho\in(0,1]$. First, due to Lemma 2.3, the series in (2.8) converges absolutely and uniformly on every compact set contained in ${\mathbb{C}}\setminus\bigcup_{m\in\Lambda}(S+m)$, hence, the function $\varphi^{(\rho)}$ is analytic in ${\mathbb{C}}\setminus\bigcup_{m\in\Lambda}(S+m)$. Since $\phi$ is meromorphic and the poles are contained in $Q$, it follows that $\varphi^{(\rho)}$ is meromorphic in $\mathbb{C}$}. Also, since $\varphi$ is meromorphic, $\psi^{(\rho)}$ is meromorphic in $\mathbb{C}$}, too.
[…]
(2.12) $\varphi^{(\rho)}(z)=\sum_{n=1}^{N}\frac{a_{n}^{(\rho)}}{(z-z_{0})^{n}}+\alpha^{(\rho)}(z)$
for some complex numbers $a_{n}^{(\rho)}$ and some function $\alpha^{(\rho)}$ which is analytic in a neighborhood $U$ of $z_{0}$.
If $z_{0}\in\partial(Q+m_{0}))$, we can choose $U$ so small that $|z-m|\geq 1/2$ for all $z\in U$ with $|m-m_{0}|\geq 2$. By (2.3), $|\phi(z-m)|\leq C_{\phi}(|m-m_{0}|+1)^{-b}$ for all such $m$ and all $z\in U$, hence,
(2.13) $\sum_{|m-m_{0}|\geq 2}2^{-\rho|m|}|\phi(z-m)|\leq\sum_{|m-m_{0}|\geq 2}C_{\phi}(|m-m_{0}|+1)^{-b}\leq C_{\phi}^{\prime}$
for $z\in U$. We conclude that the pole of $\varphi^{(\rho)}$ at $z_{0}$ is determined only by the terms with the functions $\phi(z-m)$ with $|m-m_{0}|\leq 1$ in the series (2.8). But this is not possible, since the poles of $\phi$ were assumed to be in (the interior of) $Q$, not on the boundary, see condition $({\mathcal{A}})$. Hence, $z_{0}\notin\partial(Q+m_{0})$.
We can thus choose $U$ such that dist $(U,\partial(Q+m_{0}))\geq\delta$ for some $\delta>0$, and we obtain that $|z-m|\geq\delta$ for all $z\in U$, $m\not=m_{0}$. Again, $|\phi(z-m)|\leq C(|m-m_{0}|+1)^{-b}$ for some constant depending on $\phi$ only and for all $z\in U$ and $m\not=m_{0}$, which yields
(2.14) $\sum_{m\not=m_{0}}2^{-\rho|m|}|\phi(z-m)|\leq\sum_{m\not=m_{0}}|\phi(z-m)|\leq\sum_{m\not=m_{0}}C(|m-m_{0}|+1)^{-b}\leq C_{\phi}^{\prime}$
for some positive constant $C_{\phi}^{\prime}$, independent of $\rho$ (we will need this later). The pole of $\varphi^{(\rho)}$ at $z_{0}$ is thus determined only by the term with the function $\phi(z-m_{0})$ in the series (2.8). More precisely, we deduce by comparing (2.8), (2.12) and (2.14) that
(2.15) $\phi(z-m_{0})=\sum_{n=1}^{N}\frac{2^{\rho|m_{0}|}a_{n}^{(\rho)}}{(z-z_{0})^{n}}+\beta^{(\rho)}(z)$
for some function $\beta^{(\rho)}$ which is analytic in a neighborhood of $z_{0}$.
Due to (2.3) and an estimation similar to (2.6), the bound (2.14) holds also in the case $\rho=0$, which, in view of (2.2) and (2.15) implies that also
(2.16) $\varphi(z)=\sum_{n=1}^{N}\frac{2^{\rho|m_{0}|}a_{n}^{(\rho)}}{(z-z_{0})^{n}}+\gamma^{(\rho)}(z)$
for another function $\gamma^{(\rho)}$ analytic in a neighborhood of $U$.
We conclude that in the definition of $\psi^{(\rho)}$, the poles of $\varphi^{(\rho)}$ and $\varphi$ at the point $z_{0}$ cancel each other, i.e., the function $\psi^{(\rho)}$ is analytic in $U$. Since $z_{0}$ was an arbitrary pole of $\psi^{(\rho)}$, the function $\psi^{(\rho)}$ must be analytic in $\Omega$.
$(ii)$ We first claim that, for any $m\in\Lambda$,
(2.17) $z\mapsto\frac{\phi(z-m)}{\varphi(z)}$
is a bounded function in $\Omega$. Namely, by formulas (2.15) and (2.16), the function $\phi(z-m)/\varphi(z)$ is analytic in a neighborhood of the finitely many poles of $\phi(z-m)$. Outside this neighborhood, the function $\phi(z-m)$ is uniformly bounded, due to the assumption (2.3). Also, by Remark 2.2, $|\varphi(z)|\geq C$ for a positive constant $C$, for all $z\in\Omega$. Putting these observations together proves the claim.
We next show that (2.9) holds. Namely, if $m_{0}$ is such that $z\in{\rm cl}(\varpi_{m_{0}})$, then we have
(2.18) $|\psi^{(\rho)}(z)|\leq\frac{|\phi(z-m_{0})|}{|\varphi(z)|}+\frac{1}{|\varphi(z)|}\sum_{\stackrel{{\scriptstyle\scriptstyle m\in\Lambda}}{{m\not=m_{0}}}}|\phi(z-m)|,$
and here, the first term has an upper bound independent of $\rho$ or $z$ by the remark above. Also, the second term is bounded a constant independent of $\rho$ or $z$, by Remark 2.2, (2.3) and an argument similar to (2.14). We obtain (2.9).
To prove the uniform convergence, let $M\in{\mathbb{N}}$ and $\varepsilon>0$ be given. Denoting $K=\big([-M,M]\times[-M,M]\big)\cap\Omega$, we use (2.3) (see also the calculation in (2.14)) to choose $N\in{\mathbb{N}}$ such that
(2.19) $\sup_{z\in K}\sum_{\stackrel{{\scriptstyle\scriptstyle m\in\Lambda}}{{|m|>N}}}|\phi(z-m)|\leq\sum_{\stackrel{{\scriptstyle\scriptstyle m\in\Lambda}}{{|m|>N}}}\sup_{z\in K}|\phi(z-m)|<\varepsilon.$
Then, by (2.2),
$|\psi^{(\rho)}-1|=\frac{1}{|\varphi(z)|}\Big|\sum_{m\in\Lambda}2^{-\rho|m|}\phi(z-m)-\sum_{m\in\Lambda}\phi(z-m)\Big|$
$\leq$ $\sum_{\stackrel{{\scriptstyle\scriptstyle m\in\Lambda}}{{|m|\leq N}}}\big(1-2^{-\rho|m|}\big)\frac{|\phi(z-m)|}{|\varphi(z)|}$
$\ \ +$ $\frac{1}{|\varphi(z)|}\Big|\sum_{\stackrel{{\scriptstyle\scriptstyle m\in\Lambda}}{{|m|>N}}}2^{-\rho|m|}\phi(z-m)\Big|+\frac{1}{|\varphi(z)|}\Big|\sum_{\stackrel{{\scriptstyle\scriptstyle m\in\Lambda}}{{|m|>N}}}\phi(z-m)\Big|.$
Here, since $|\varphi(z)|\geq C>0$ for $z\in\Omega$, the last two terms are bounded by $\varepsilon/C$, by (2.19). Also, the first term can be estimated using the boundedness of the finitely many expressions (2.17),
$\sum_{\stackrel{{\scriptstyle\scriptstyle m\in\Lambda}}{{|m|\leq N}}}\big(1-2^{-\rho|m|}\big)\frac{|\phi(z-m)|}{|\varphi(z)|}\leq C_{N}\big(1-2^{-\rho N}\big),$
which is bounded by $\varepsilon$, if $\rho$ is small enough. Hence, $\psi^{(\rho)}\to 1$ uniformly on $\big([-M,M]\times[-M,M]\big)\cap\Omega$.
$(iii)$ If $f\in L^{2}(\Omega)$ and $\varepsilon>0$ is given, we pick up $M\in{\mathbb{N}}$ so large that
(2.20) $\int\limits_{\Omega\cap\{|z|\geq M\}}\!\!\!\!\!\!\!\!(1+C)^{2}|f|^{2}dA<\varepsilon,$
where $C$ is the constant in (2.9). Then, by (2.20),
(2.21) $\int\limits_{\Omega}|f\psi^{(\rho)}-f|^{2}dA\leq\!\!\!\!\!\!\int\limits_{\Omega\cap\{|z|\leq M\}}\!\!\!\!\!\!|1-\psi^{(\rho)}|\,|f|^{2}dA+\!\!\!\!\!\!\int\limits_{\Omega\cap\{|z|\geq M\}}\!\!\!\!\!\!(1+C)^{2}|f|^{2}dA$
$\leq$ $\|f\|_{\Omega}^{2}\sup_{|z|\leq M}|1-\psi^{(\rho)}(z)|+\varepsilon.$
The first term on the right hand side can be made smaller than $\varepsilon$, if $\rho$ is small enough, by using the already proven claim $(ii)$ of the theorem, which proves the claim $(iii)$.
$(iv)$ Let $K$ and $m$ be given and $z\in K$. We have
$|\psi^{(\rho)}(z+m)|\leq\!\!\!\!\!\!\sum_{\stackrel{{\scriptstyle\scriptstyle n\in\Lambda}}{{|n-m|\leq|m|/2}}}\!\!\!\!\!\!2^{-\rho|n|}\frac{|\phi(z+m-n)|}{|\varphi(z)|}+\!\!\!\!\!\!\sum_{\stackrel{{\scriptstyle\scriptstyle n\in\Lambda}}{{|n-m|>|m|/2}}}\!\!\!\!\!\!2^{-\rho|n|}\frac{|\phi(z+m-n)|}{|\varphi(z)|}$
$=:R_{1}+R_{2}.$
To estimate $R_{1}$, the summation index satisfies $|n|\geq|m|/2$, hence, $2^{-\rho|n|}\leq 2^{-\rho|m|/2}$ and this yields
$R_{1}\leq 2^{-\rho|m|/2}\sum_{n\in\Lambda}\frac{|\phi(z+m-n)|}{|\varphi(z)|}.$
Here, the sum is bounded by a constant independent of $z,m$, by an argument similar to that following (2.18). Hence,
$R_{1}\leq C2^{-\rho|m|/2}\leq\frac{C^{\prime}}{\rho^{b}(1+|m|)^{b}}.$
As for $R_{2}$,
(2.22) $R_{2}\leq\!\!\!\!\!\!\sum_{\stackrel{{\scriptstyle\scriptstyle n\in\Lambda}}{{|n-m|>|m|/2}}}\!\!\!\!\!\!2^{-\rho|n|}|m-n|^{-b}\frac{|m-n|^{b}|\phi(z+m-n)|}{|\varphi(z)|}$
$\leq$ $|m/2|^{-b}\sup_{w\in K,n\in\Lambda}\frac{|n|^{b}|\phi(w+n)|}{|\varphi(z)|}\sum_{k\in\Lambda}2^{-\rho|k|}.$
Let $M\in{\mathbb{N}}$ be so large that $K\subset[-M+2,M-2]\times[-M+2,M-2]$. Then, $|n|^{b}\leq C_{K}|n+w|^{b}$ for all $w\in K$ and $n\in\Lambda$ with $|n|\geq 2M$, hence, (2.17), (2.3) and Remark 2.2 imply
(2.23) $\sup_{w\in K,n\in\Lambda}\!\!\!\!\frac{|n|^{b}|\phi(w+n)|}{|\varphi(z)|}$
$\leq$ $C_{K}\!\!\!\!\sup_{w\in K,|n|<2M}\!\!\!\!\frac{|\phi(w+n)|}{|\varphi(z)|}+C_{K}\!\!\!\!\sup_{w\in K,|n|\geq 2M}\!\!\!\!|w+n|^{b}|\phi(w+n)|\leq C_{K}^{\prime}.$
Since
$\sum_{k\in\Lambda}2^{-\rho|k|}\leq\frac{C}{\rho^{2}},$
we get from (2.23) that (2.22) is also bounded by $C\rho^{-b}(1+|m|)^{-b}.$ $\Box$
We will need in the next sections an invertible analytic multiplier function for switching the Floquet parameter in the quasiperiodic boundary conditions. Such a function can be constructed with the help of a second modification of the elliptic function $\varphi$ in condition $({\mathcal{A}})$. The definition will be given in two steps. Recall that we denote $m_{1}={\rm Re}\,m$, $m_{2}=\ {\rm Im}\,m$ for all $m\in\Lambda$.
Lemma 2.5.
Let $\varphi$ and $\phi$ be as in (2.2) and $\eta\in{\bf Q}$ and define the functions
(2.24) $\varphi_{\eta}(z)=\sum_{m\in\Lambda}e^{-i\eta_{1}m_{1}-i\eta_{2}m_{2}}\phi(z-m),\ z\in\Omega,\ \ \ \mbox{and}\ \ \ {\widetilde{\psi}}_{\eta}=\frac{\varphi_{\eta}}{\varphi}.$
$(i)$ For every $\eta\in{\bf Q}$, the functions $\varphi_{\eta}$ and ${\widetilde{\psi}}_{\eta}$ are meromorphic in $\mathbb{C}$}, and ${\widetilde{\psi}}_{\eta}$ is analytic on $\Omega$.
$(ii)$ There exist constants $\beta>0$ and $C>0$ such that
(2.25) $\sup_{z\in\Omega,\eta\in{\bf Q}}|{\widetilde{\psi}}_{\eta}(z)|\leq C\ \ \ \ \mbox{and}\ \ \ \ \sup_{z\in\varpi}|{\widetilde{\psi}}_{\eta}(z)-1|\leq C|\eta|^{\beta}\ \forall\,\eta\in{\bf Q}.$
$(iii)$ The function ${\widetilde{\psi}}_{\eta}$ satisfies the quasiperiodic boundary conditions
(2.26) ${\widetilde{\psi}}_{\eta}(m+1+iy)=e^{i\eta_{1}}{\widetilde{\psi}}_{\eta}(m+iy),\ \ \ {\widetilde{\psi}}_{\eta}(m+x+i)=e^{i\eta_{2}}{\widetilde{\psi}}_{\eta}(m+x)$
for all $\eta\in{\bf Q}$, $m\in\Lambda$ and $x,y\in(0,1)$.
Proof. Assertion $(i)$ can be proved in the same way as Lemma 2.4.(i)(i), since the moduli of the factors $e^{-i\eta_{1}m_{1}-i\eta_{2}m_{2}}$ equal one. As for $(ii)$, the first bound in (2.25) can be proved in the same way as (2.9).
Let us prove the second inequality in (2.25). We have for all $z\in{\rm cl}(\varpi)\cap\Omega$ and nonzero $\eta\in{\bf Q}$, by Remark 2.2,
$|{\widetilde{\psi}}_{\eta}(z)-1|\leq\frac{\sum_{m\in\Lambda}|1-e^{-i\eta_{1}m_{1}-i\eta_{2}m_{2}}|\,|\phi(z-m)|}{\Big|\sum_{m\in\Lambda}\phi(z-m)\Big|}$
$\leq$ $\frac{\sum_{|m|\leq|\eta|^{-1/2}}|1-e^{-i\eta_{1}m_{1}-i\eta_{2}m_{2}}|\,|\phi(z-m)|}{\Big|\sum_{m\in\Lambda}\phi(z-m)\Big|}$
$\ \ +$ $C^{\prime}\!\!\!\!\!\sum_{|m|>|\eta|^{-1/2}}\!\!\!\!\!|1-e^{-i\eta_{1}m_{1}-i\eta_{2}m_{2}}|\,|\phi(z-m)|=:S_{1}+S_{2}.$
To evaluate $S_{1}$ we note that, by the Taylor series of the exponential function,
$|1-e^{-i\eta_{1}m_{1}-i\eta_{2}m_{2}}|\leq C|\eta|\,|m|\leq C|\eta|^{1/2},$
which yields
(2.27) $S_{1}\leq C\frac{|\eta|^{1/2}\sum_{m\in\Lambda}|\phi(z-m)|}{\big|\sum_{m\in\Lambda}\phi(z-m)\big|}\leq C^{\prime}|\eta|^{1/2},$
where we at the end again used an argument similar to that following (2.18).
For $S_{2}$, we use $|1-e^{-i\eta_{1}m_{1}-i\eta_{2}m_{2}}|\leq 2$ and an argument similar to (2.6) to get
$S_{2}\leq\!\!\!\!\!\sum_{|m|>|\eta|^{-1/2}}\!\!\!\!\!\,|m|^{2+\delta-b}|m|^{b-2-\delta}|\phi(z-m)|$
$\leq$ $|\eta|^{(b-2-\delta)/2}\sum_{|m|>|\eta|^{-1/2}}|m|^{b-2-\delta}|\phi(z-m)|\leq|\eta|^{(b-2-\delta)/2};$
here, it is necessary to assume that $|\eta|<c_{0}$ for some small enough constant $c_{0}>0$ so that, say, $|z-m|\geq 2$ holds for $z\in\varpi$ and $m$ with $|m|>|\eta|^{-1/2}>c_{0}^{-1/2}$. This is not a restriction, since we have already proved the first inequality in (2.25) and thus the second one automatically holds for $|\eta|\geq c_{0}$, once its constant $C$ is large enough.
$(iii)$ The defining formula (2.24) implies that $\varphi_{\eta}$ satisfies the quasiperiodic boundary conditions (2.26), hence, also $\psi_{\eta}$ satisfies them, since $1/\varphi$ is doubly periodic. $\Box$
As mentioned above, we will need an invertible multiplier operator changing the Floquet parameter values. Now, the second formula (2.25) implies that $\widetilde{\psi}_{\eta}$ is an invertible function, if $|\eta|$ is small enough. However, there does not seem be any guarantee that this is true for all $\eta\in{\bf Q}$: the function $\varphi_{\eta}$ and thus $\widetilde{\psi}_{\eta}$ could have zeros in the domain $\varpi$, which would destroy the invertibility of the associated multiplier operator.
Let $R>0$ be such that
(2.28) $\sup_{z\in\varpi}|{\widetilde{\psi}}_{\eta}(z)-1|\leq\frac{3}{4}\ \ \ \mbox{for all}\ |\eta|\leq R;$
such an $R$ can always be found, by (2.25). Clearly, (2.28) implies
(2.29) $\inf_{z\in\varpi}|\widetilde{\psi}_{\eta}(z)|\geq\frac{1}{4}\ \ \ \mbox{for all}\ |\eta|\leq R.$
In the case $R\geq\sqrt{2}\pi$, the inequality (2.29) holds for all $\eta\in{\bf Q}$, and we define $\psi_{\eta}=\widetilde{\psi}_{\eta}$ for all $\eta\in{\bf Q}$. If $R<\sqrt{2}\pi$, we define
(2.30) $\psi_{\eta}(z)=\left\{\begin{array}[]{ll}\widetilde{\psi}_{\eta}(z),&\mbox{for $\eta\in{\bf Q}$ with }|\eta|\leq R\\ \widetilde{\psi}_{R\eta/|\eta|}(z)^{|\eta|/R},&\mbox{for $\eta\in{\bf Q}$ with }|\eta|>R.\end{array}\right$.
If $z\in\varpi$, then $\widetilde{\psi}_{R\eta/|\eta|}(z)=re^{i\theta}$ with $\theta\in(-\pi/2,\pi/2)$, and the power in (2.30) is defined as
$\widetilde{\psi}_{R\eta/|\eta|}(z)^{|\eta|/R}=r^{|\eta|/R}e^{i\theta|\eta|/R}.$
Lemma 2.6.
Let $\eta\in{\bf Q}$.
$(i)$ The function $\psi_{\eta}$ is analytic on $\Omega$, and there exists a constant $C>0$ such that
(2.31) $\sup_{z\in\Omega,\eta\in{\bf Q}}|\psi_{\eta}(z)|\leq C.$
$(ii)$ If $\beta>0$ is as in (2.25), there holds
(2.32) $\sup_{z\in\varpi}|1-\psi_{\eta}(z)|\leq\min\big(3/4,C|\eta|^{\beta}\big)\ \ \ \mbox{and}\ \ \ \inf_{z\in\varpi,\eta\in{\bf Q}}|\psi_{\eta}(z)|\geq 1/4.$
$(iii)$ The function $\psi_{\eta}$ satisfies the quasiperiodic boundary conditions
(2.33) $\psi_{\eta}(m+1+iy)=e^{i\eta_{1}}\psi_{\eta}(m+iy),\ \ \ \psi_{\eta}(m+x+i)=e^{i\eta_{2}}\psi_{\eta}(m+x)$
for all $m\in\Lambda$, $x,y\in(0,1)$.
Proof. Given $\eta$, $\psi_{\eta}$ is analytic in $\Omega$ due to its definition (2.30), since all positive real powers of $\widetilde{\psi}_{\eta}$ can be defined as analytic functions, see the second inequality in (2.25). All other claims in $(i)$ and $(ii)$ follow from Lemma 2.5, (2.28), (2.29) and the definitions after them. Note that the upper bound $C|\eta|^{\beta}$ in (2.32) only concerns $\eta$ in some neighborhood of 0, where $\psi_{\eta}$ and $\widetilde{\psi}_{\eta}$ coincide, hence (2.32) follows from (2.25).
The quasiperiodicity (2.33) follows from (2.26), if $|\eta|\leq R$. If $|\eta|>R$, we have
$\psi_{\eta}(m+1+iy)={\widetilde{\psi}}_{R\eta/|\eta|}(m+1+iy)^{|\eta|/R}$
$=$ $\big(e^{iR\eta_{1}/|\eta|}\big)^{|\eta|/R}{\widetilde{\psi}}_{R\eta/|\eta|}(m+iy)^{|\eta|/R}=e^{i\eta_{1}}\psi_{\eta}(m+iy),$
for all $m\in\Lambda$ and $y\in(0,1)$ and, similarly, $\psi_{\eta}(m+x+i)=e^{i\eta_{2}}\psi_{\eta}(m+x)y$ for all $x\in(0,1)$. $\Box$
3. Mapping properties of the Floquet transform in Bergman spaces.
We proceed to the first main result of the paper, namely, the characterization of the image ${\sf F}(A^{2}(\Omega))\subset L^{2}({\bf Q};L^{2}(\varpi))$ in Theorem 3.4. Naturally, this result is of importance, since one wants to preserve the bijectivity properties of $\sf F$} also in the Bergman space case. Based on Theorem 3.6 in [10], one cannot expect that just replacing $L^{2}(\varpi)$ by $A^{2}(\varpi)$ in the Bochner space makes the restriction of $\sf F$} to $A(\Omega)$ a surjection: the quasiperiodic boundary conditions should appear in ${\sf F}(A^{2}(\Omega))$. In this section we show that with the help of the results in Section 2, the approach of [10] also works in the doubly periodic case. We will omit some details and refer to [10] for them.
In fact, the range of the ${\sf F}(A^{2}(\Omega))$ appears in the next definition. It involves the subspace $A_{\eta,{\rm ext}}^{2}(\varpi)$ consisting of Bergman functions satisfying quasiperiodic boundary conditions. Formula (3.2) will give a lot of examples of such functions.
Definition 3.1.
If $\eta\in{\bf Q}$, we denote by $A_{\eta,{\rm ext}}^{2}(\varpi)$ the subspace of $A^{2}(\varpi)$ consisting of functions $f$ which can be extended as analytic functions in a neighborhood of ${\rm cl}(\varpi)\cap\Omega$ in $\Omega$ and satisfy the quasiperiodic boundary conditions
(3.1) $f(1+iy)=e^{i\eta_{1}}f(iy),\ f(x+i)=e^{i\eta_{2}}f(x)\ \ \ \mbox{for all}\ x,y\in(0,1).$
Moreover, $A_{\eta}^{2}(\varpi)$ stands for the closure of $A_{\eta,{\rm ext}}^{2}(\varpi)$ in $A^{2}(\varpi)$ and $L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$ for the subspace of $L^{2}({\bf Q};A^{2}(\varpi))$ consisting of functions $f$ such that the function $z\mapsto f(z,\eta)$ belongs to $A_{\eta}^{2}(\varpi)$ for a.e. $\eta\in{\bf Q}$.
We continue with the following observations.
Proposition 3.2.
Assume condition $({\mathcal{A}})$ holds for $\Omega$.
$(i)$ For every $f\in A^{2}(\Omega)$, the function $z\mapsto{\sf F}f(z,\eta)$ is analytic on $\varpi$ for a.e. $\eta\in{\bf Q}$.
$(ii)$ The space $A^{2}(\Omega)$ has a dense subspace, denoted by $X$, which consists of functions $f$ such that, for all $\eta\in{\bf Q}$, the function ${\sf F}f(\cdot,\eta)$ belongs to $A_{\eta,{\rm ext}}^{2}(\varpi)$.
Proof.(i)(i) Let $\rho\in(0,1]$ and let $\psi^{(\rho)}$ be as in Lemma 2.4. We first show that for every $f\in A^{2}(\Omega)$ and $\eta\in{\bf Q}$, the function
(3.2) ${\sf F}(\psi^{(\rho)}f)(z,\eta)=\frac{1}{2\pi}\sum_{m\in\Lambda}e^{-i\eta m}\psi^{(\rho)}(z+m)f(z+m)$
is analytic in a neighborhood of $\varpi$. To this end, we define the set $U\subset\Omega$ to be the interior of the closure in $\Omega$ of the set
${\textstyle\bigcup}_{|m|\leq 2}\varpi_{m}$
(note that the set $\varpi=\varpi_{0}$ is included in $U$). Then, let $\xi\in U$ be arbitrary and pick up a small enough $\varrho>0$ such that cl $\big(B(\xi,\varrho)\big)\subset U.$ Applying a translation, the Cauchy integral formula implies the estimate
$\sup\limits_{z\in B(\xi,\varrho)}|f(z+m)|=\sup\limits_{z\in B(\xi,\varrho)+m}|f(z)|$ $\leq$ $\frac{C}{\varrho}\|f\|_{\varpi+m}\leq\frac{C}{\varrho}\|f\|_{\Omega}$
for all $m\in\Lambda$ and $z\in B(\xi,\varrho)$. The estimate (2.11) implies that, for a fixed $\rho$, the series (3.2) converges uniformly in the disc $B(\xi,\varrho)$. This means, $F(\psi^{(\rho)}f)$ is analytic in $B(\xi,\varrho)$ and thus in $U\supset\varpi$.
The rest of the proof is similar to that of Proposition 3.2 of [10], except that we use (2.10) instead of (3.8) of the reference. Indeed, (2.10) and the unitarity of $\sf F$} imply that ${\sf F}(\psi^{(\rho)}f)$ converges to ${\sf F}f$ in $L^{2}({\bf Q};L^{2}(\varpi))$ as $\rho\to 0$, and we thus find a decreasing sequence $(\rho_{k})_{k=1}^{\infty}$ with $0<\rho_{k}\to 0$ such that
$\lim_{k\to\infty}{\sf F}(\psi^{(\rho_{k})}f)(\cdot,\eta)=Ff(\cdot,\eta)\ \ \mbox{in}\ L^{2}(\varpi)$
for almost all $\eta\in{\bf Q}$. This implies that, for these $\eta$, the function ${\sf F}f$ is analytic in $\varpi$, since ${\sf F}(\psi^{(\rho)}f)(\cdot,\eta)$ is analytic in $\varpi$.
$(ii)$ The dense subspace $X$ can be defined to consist of all functions $\psi^{(\rho)}f$, where $f\in A^{2}(\Omega)$. As it was shown above, every ${\sf F}(\psi^{(\rho)}f)$ is an analytic function in $U$, which contains a neighborhood of cl $(\varpi)$. It follows from the defining formula (3.2) that every ${\sf F}(\psi^{(\rho)}f)$ satisfies the quasiperiodic conditions (3.1). Finally, the density of $X$ in $L^{2}(\Omega)$ follows from Lemma 2.4.(iii)(iii). $\Box$
We still consider some remarks and definitions, which are needed in proof of Theorem 3.4. First, we observe that the subspace $L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$ of $L^{2}({\bf Q};L^{2}(\varpi))$ is closed, since $A_{\eta}^{2}(\varpi)$ is closed in $A^{2}(\varpi)$ for every $\eta$. We also define
${\mathcal{H}}_{\eta}:=L^{2}\big({\bf Q};A_{\eta,{\rm ext}}^{2}(\varpi)\big)$
to be the subspace of $L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$ which consists of functions $g$ such that the mapping $z\mapsto g(z,\eta)$ belongs to $A_{\eta,{\rm ext}}^{2}(\varpi)$ for a.e. $\eta$.
The following lemma can be proved in the same way as Lemma 3.5 of [10], except that one uses the function $\psi_{\eta}$ of Lemma 2.6.(iii)(iii) instead of $e^{i\eta z}$ of the reference. Note that the proof is not completely trivial, since the spaces $L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$ and ${\mathcal{H}}_{\eta}$ only have the structure of a Banach vector bundle (e.g. Section 1.3 of [5] or [10], p. 209) and simple functions of $L^{2}\big({\bf Q};A^{2}(\varpi)\big)$ are not in general contained in these subspaces.
Lemma $3.$
The space ${\mathcal{H}}_{\eta}$ is a dense subspace of $L^{2}\big({\bf Q};A_{\eta}^{2}(\varpi)\big)$.
With these preparations we can state and prove the main result of this section.
Theorem 3.4.
Let $\Omega$ be doubly periodic domain satisfying assumption $({\mathcal{A}})$. The Floquet transform $\sf F$} is a unitary operator from $A^{2}(\Omega)$ onto $L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$ with inverse ${\sf F}^{-1}:L^{2}({\bf Q};A_{\eta}^{2}(\varpi))\to A^{2}(\Omega)$ given by the formula (1.3).
Proof. We first observe that by Proposition 3.2.(ii)(ii), $\sf F$} maps the dense subspace $X$ of $A^{2}(\Omega)$ into ${\mathcal{H}}_{\eta}$. In view of the unitarity of $\sf F$}, this implies that ${\sf F}(A^{2}(\Omega))$ is contained in the closure of ${\mathcal{H}}_{\eta}$ in $L^{2}({\bf Q};L^{2}(\varpi))$, which equals $L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$, by Lemma $3.$
We are thus left with the construction of an inverse image under $\sf F$} for all $g\in L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$. Due to the density of ${\mathcal{H}}_{\eta}$ in $L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$ and the unitarity of $\sf F$}, we may assume that $g\in{\mathcal{H}}_{\eta}$. In the next construction we consider $\eta\in{\bf Q}$ such that the function $g(\cdot,\eta)$ is a well-defined element of $A_{\eta,{\rm ext}}^{2}(\varpi)$. For all such $\eta\in{\bf Q}$, we define the function $G_{\eta}:\Omega\to{\mathbb{C}}$ by setting
(3.3) $G_{\eta}(z)=G_{\eta,m}(z)\ \ \ \mbox{for all}\ z\in{\rm cl}(\varpi_{m}),\ m\in\Lambda,$
where
(3.4) $G_{\eta,m}:\varpi_{m}\to{\mathbb{C}},\ \ G_{\eta,m}(z)=e^{i\eta_{1}m_{1}+i\eta_{2}m_{2}}g(z-m,\eta).$
Let us show that $G_{\eta}$ is well defined and gives an analytic function in $\Omega$. Indeed, $G_{\eta,m}$ is analytic on the subdomain $\varpi_{m}$, $m\in\Lambda$ (see (1.1)) and has an analytic extension, still denoted by $G_{\eta,m}$, to a neighborhood of the closure of $\varpi_{m}$ in $\Omega$ (see Definition 3.1). Moreover, the functions $g(\cdot,\eta)$ satisfy the quasiperiodic boundary conditions (3.1). Hence, we get for all $m\in\Lambda$, all $z=m+1+iy$ with $y\in(0,1)$,
$G_{\eta,m}(z)=e^{i\eta_{2}m_{2}}e^{i\eta_{1}m_{1}}g(z-m,\eta)=e^{i\eta_{2}m_{2}}e^{i\eta_{1}m_{1}}g(1+iy,\eta)$
$=$ $e^{i\eta_{2}m_{2}}e^{i(m_{1}+1)\eta_{1}}g(iy,\eta)=e^{i\eta_{2}m_{2}}e^{i(m_{1}+1)\eta_{1}}g(z-(m+1),\eta)=G_{\eta,m+1}(z).$
We see that the functions $G_{\eta,m}$ and $G_{\eta,m+1}$ coincide on the line segments of positive length, which consist of the common parts of $\partial\varpi_{m}$ and $\partial\varpi_{m+1}$. Due to the their analyticity, $G_{\eta,m}$ and $G_{\eta,m+1}$ thus coincide in their entire common domain of definition. In the same way one shows that the functions $G_{\eta,m}$ and $G_{\eta,m+i}$ coincide in their (non-empty) common domain of definition. Since these claims hold for all $m\in\Lambda$, we conclude that the function $G_{\eta}$ in (3.3) is well-defined and analytic in $\Omega$.
From (3.4), (1.3) we observe that
$H:=\frac{1}{2\pi}\int\limits_{{\bf Q}}G_{\eta}(z)d\eta={\sf F}^{-1}g(z)\in A^{2}(\Omega),$
thus, there holds ${\sf F}H=g$ for all $g\in{\mathcal{H}}_{\eta}\subset L^{2}\big({\bf Q};L^{2}(\varpi)\big)$. We conclude that image of ${\sf F}\big(A^{2}(\Omega)\big)$ contains the subspace ${\mathcal{H}}_{\eta}$. In view of the fact that $\sf F$} is an isometry, ${\sf F}\big(A^{2}(\Omega)\big)$ also contains the entire space $L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$. This completes the proof. $\Box$
4. General kernel formula for the periodic domain.
The results of the previous section identify the $\eta$-dependent family of Bergman spaces $A_{\eta}^{2}(\varpi)$ on the periodic cell $\varpi$, which corresponds, under the Floquet transform, to the Bergman space on the original doubly periodic domain $\Omega$. The quasiperiodic boundary conditions in the periodic cell do not show up in the corresponding $L^{2}$-isometry (see the beginning of Section 1), but the situation in the Bergman space case is analogous to the Sobolev space case, see [6], Section 4.2. Let us next present the connection of the Bergman-type projections in the domains $\Omega$ and $\varpi$ by using the Floquet transform. The results and their proofs are completely analogous to the singly periodic case so that a repetition of the proofs will not be necessary.
If $\eta\in{\bf Q}$, let us denote the orthogonal projection from $L^{2}(\varpi)$ onto $A_{\eta}^{2}(\varpi)$ by $P_{\eta}$. The projection $P_{\eta}$ can be written as an integral operator
(4.1) $P_{\eta}f(z)=\int\limits_{\varpi}K_{\eta}(z,w)f(w)dA(w),$
since the existence of kernel can be proved following the usual proof for the existence of Bergman kernels, see [4], p.1060. In particular, the kernel has the property that $K_{\eta}(z,\cdot)\in L^{2}(\varpi)$ for all $z\in\varpi$.
We next define the bounded operator ${\mathcal{P}}:L^{2}({\bf Q};L^{2}(\varpi))\to L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$ by
(4.2) ${\mathcal{P}}f(z,\eta)=\big(P_{\eta}f(\cdot,\eta)\big)(z),\ \ \ f\in L^{2}({\bf Q};L^{2}(\varpi)),\ z\in\varpi,\ \eta\in{\bf Q}.$
The following result can now be proved in the same way as Lemma 4.2 and Theorem 4.3 in [10]. We denote here $z^{\mathsf{tr}}=z-[{\rm Re}z]-i[{\rm Im}z]$ so that $z^{\mathsf{tr}}\in{\rm cl}(\varpi)$ for all $z\in\Omega$.
Theorem 4.1.
Let $\Omega$ be doubly periodic domain satisfying assumption $({\mathcal{A}})$.
$(i)$ The operator $\mathcal{P}$} is the orthogonal projection from $L^{2}({\bf Q};L^{2}(\varpi))$ onto $L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$.
$(ii)$ The Bergman projection $P_{\Omega}$ from $L^{2}(\Omega)$ onto $A^{2}(\Omega)$ can be written as
(4.3) ${\sf F}^{-1}{\mathcal{P}}{\sf F}f(z)=\frac{1}{2\pi}\int\limits_{\bf Q}e^{i[{\rm Re}z]\eta_{1}+i[{\rm Im}z]\eta_{2}}\big(P_{\eta}\widehat{f}(\cdot,\eta)\big)(z^{\mathsf{tr}})d\eta$
$=$ $\frac{1}{4\pi^{2}}\int\limits_{\Omega}\int\limits_{\bf Q}e^{i\eta_{1}([{\rm Re}z]-[{\rm Re}w])+i\eta_{2}([{\rm Im}z]-[{\rm Im}w])}K_{\eta}(z^{\mathsf{tr}},w^{\mathsf{tr}})f(w)d\eta dA(w)$
5. Spectral band formula for Toeplitz operators with doubly periodic symbols
Given a doubly periodic domain $\Omega$ and a symbol $a\in L^{\infty}(\Omega)$, the Toeplitz operator $T_{a}:A^{2}(\Omega)\to A^{2}(\Omega)$ is defined by
$T_{a}f=P_{\Omega}M_{a}f=P_{\Omega}(af)$
where $M_{a}$ is the pointwise multiplier $f\mapsto af$, $f\in A^{2}(\Omega)$. The aim of this section is to study the spectra $\sigma(T_{a})$ and essential spectra $\sigma_{\rm ess}(T_{a})$ of the operators $T_{a}$ with doubly periodic symbols $a\in L^{\infty}(\Omega)$,
(5.1) $a(z)=a(z+1)=a(z+i)\text{ for almost all }z\in\Omega.$
In the main result, Theorem 5.1, we show that the spectrum and essential spectrum of such a $T_{a}$ coincide and, moreover, can be presented as the union of the spectra of a family of Toeplitz-type operators $T_{a,\eta}:A_{\eta}^{2}(\omega)\to A_{\eta}^{2}(\omega)$, $\eta\in{\bf Q}$, in the periodic cell. These operators are defined with the help of the projections $P_{\eta}$ of (4.1) by
$T_{a,\eta}f=P_{\eta}(a|_{\varpi}f)$
for all $\eta\in{\bf Q}$, $f\in A_{\eta}^{2}(\omega)$, and their spectra as operators in $A_{\eta}^{2}(\omega)$ are denoted by $\sigma(T_{a,\eta})$.
The main result reads as follows.
Theorem 5.1.
Let $\Omega$ be doubly periodic domain with property $({\mathcal{A}})$ and let the symbol $a\in L^{\infty}(\Omega)$ be as in (5.1). The essential spectrum of the Toeplitz-operator $T_{a}:A^{2}(\Omega)\to A^{2}(\Omega)$ can be described by the formula
(5.2) $\sigma_{\rm ess}(T_{a})=\bigcup_{\eta\in{\bf Q}}\sigma(T_{a,\eta}).$
Moreover, there holds $\sigma(T_{a})=\sigma_{\rm ess}(T_{a})$.
The proof follows that of Theorem 4.1. of [11], but the main difference is the need to use the multiplier $J_{\mu,\eta}$, the definition of which is based on the much more delicate function $\psi_{\eta}$ of Lemma 2.6 instead of the basic exponential function $e^{i\eta z}$ of [11]. This causes some differences to the arguments. For example, the function $1/\psi_{\eta}$ is not the same as $\psi_{-\eta}$ as in the case of the exponential function, which could complicate the considerations involving the inverse operators, and the estimate (5.4) in below is worse than its analogue (3.9) in [11]. It is thus worthwhile to present the main steps of the proof.
We extend the definition of the operators $T_{a,\eta}$ to the Bochner space by defining $\mathcal{T}_{a}:L^{2}({\bf Q};A_{\eta}^{2}(\varpi))\to L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$ as
$\mathcal{T}_{a}:f(\cdot,\eta)\mapsto T_{a,\eta}f(\cdot,\eta).$
Also, we denote by $\mathcal{M}_{a}:L^{2}({\bf Q};A_{\eta}^{2}(\varpi))\to L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$ the operator
$\mathcal{M}_{a}:f(\cdot,\eta)\mapsto a|_{\varpi}f(\cdot,\eta)$
so that there holds $\mathcal{T}_{a}=\mathcal{P}\mathcal{M}_{a}$. The definitions easily imply $T_{a}f={\sf F}^{-1}\mathcal{T}_{a}{\sf F}f$ for all $f\in A^{2}(\Omega)$.
We assume that $({\mathcal{A}})$ holds for $\Omega$ and refer to Lemma 2.6 for the definition of the analytic function $\psi_{\eta}$. Given $\eta,\mu\in{\bf Q}$, we define the operators
(5.3) $J_{\eta,\mu}f(z)=\psi_{\mu-\eta}(z)f(z)\ \ \ \mbox{and}\ \ \ T_{a,\eta,\mu}=J_{\eta,\mu}^{-1}T_{a,\mu}J_{\eta,\mu}.$
Lemma 5.2.
Given $\eta,\mu\in{\bf Q}$, the operator $J_{\eta,\mu}$ is a bounded bijection of $A^{2}(\varpi)$ onto itself and also a bounded bijection from $A_{\eta}^{2}(\varpi)$ onto $A_{\mu}^{2}(\varpi)$. Consequently, $T_{a,\eta,\mu}$ is a bounded operator from $A_{\eta}^{2}(\varpi)$ into itself and from $A^{2}(\varpi)$ into itself.
Proof. It follows from Lemma 2.6.(i),(ii)(i),(ii) that $J_{\eta,\mu}$ is a bounded bijection of $A^{2}(\varpi)$ onto itself, and its inverse is, of course, the multiplication operator $J_{\eta,\mu}^{-1}:f\mapsto f/\psi_{\mu-\eta}$. Also, by Lemma 2.6.(iii)(iii), if $f\in A_{\eta,{\rm ext}}^{2}(\varpi)$, then the function $\psi_{\mu-\eta}f$ satisfies the quasiperiodic conditions (3.1) with the Floquet parameter $\mu\in{\bf Q}$. Thus, $J_{\eta,\mu}$ maps $A_{\eta}^{2}(\varpi)$ boundedly into $A_{\mu}^{2}(\varpi)$. Also, $J_{\eta,\mu}^{-1}$ maps $A_{\mu}^{2}(\varpi)$ into $A_{\eta}^{2}(\varpi)$, since the function $1/\psi_{\mu-\eta}$ satisfies the quasiperiodic conditions with the parameter $\eta-\mu$. We conclude that $J_{\eta,\mu}$ is also a bounded bijection from $A_{\eta}^{2}(\varpi)$ onto $A_{\mu}^{2}(\varpi)$. $\Box$
The operators $J_{\eta,\mu}$ are not isometries, since the factor $\psi_{\eta}$ is not unimodular. However, (2.32) implies for all $f\in L^{2}(\varpi)$,
$\int\limits_{\varpi}|f-\psi_{\eta}f|^{2}dA=\int\limits_{\varpi}|1-\psi_{\eta}|^{2}|f|^{2}dA\leq C|\eta|^{2\beta}\|f\|_{\varpi}^{2},$
hence,
(5.4) $\left\|I_{\varpi}-J_{\eta,\mu}\right\|_{L^{2}(\varpi)\rightarrow L^{2}(\varpi)}\leq C|\eta-\mu|^{\beta}\ \ \ \forall\mu,\eta\in{\bf Q}$
where $I_{\varpi}$ is the identity operator on $L^{2}(\varpi)$. Also, Lemma 2.6.(i),(ii)(i),(ii) implies
$\sup_{z\in\varpi}\Big|1-\frac{1}{\psi_{\eta}(z)}\Big|=\sup_{z\in\varpi}\frac{|1-\psi_{\eta}(z)|}{|\psi_{\eta}(z)|}\leq\frac{\sup_{z\in\varpi}|1-\psi_{\eta}(z)|}{\inf_{z\in\varpi}|\psi_{\eta}(z)|}\leq C|\eta|^{\beta}$
and consequently,
(5.5) $\left\|I_{\varpi}-J_{\eta,\mu}^{-1}\right\|_{L^{2}(\varpi)\rightarrow L^{2}(\varpi)}\leq C|\eta-\mu|^{\beta}\ \ \ \forall\mu,\eta\in{\bf Q}.$
We also get, by the definition of $T_{a,\eta,\mu}$,
$\|T_{a,\mu}-T_{a,\eta,\mu}\|_{A^{2}(\varpi)\to A^{2}(\varpi)}\leq C|\mu-\eta|^{\beta}.$
Lemma 5.3.
There exist constants $C,C^{\prime}>0$ such that if $\eta,\mu\in{\bf Q}$, we have $\|P_{\eta}-P_{\mu}\|_{L^{2}(\varpi)\to L^{2}(\varpi)}\leq C|\eta-\mu|^{\beta/2}$ and consequently $\|T_{a,\eta}-T_{a,\mu}\|_{A^{2}(\varpi)\to A^{2}(\varpi)}\leq C^{\prime}|\eta-\mu|^{\beta/2}$.
Proof. If $\mu\in{\bf Q}$ is fixed, the idea of the proof, according to [10], is to consider the non-orthogonal projections $\widetilde{P}_{\mu}=J_{\mu,\eta}^{-1}P_{\eta}J_{\mu,\eta}$ from $L^{2}(\varpi)$ onto $A_{\mu}^{2}(\varpi)$, where $\eta\in{\bf Q}$. For every $f\in L^{2}(\varpi)$ we denote $_{A}=P_{\mu}f, f=f$$^{\perp}=f-f_{A}$, and observe that $(P_{\mu}-{\widetilde{P}_{\mu}})f=(P_{\mu}-{\widetilde{P}_{\mu}})f^{\perp}$, since both $P_{\mu}$ and $\widetilde{P}_{\mu}$} project onto $A_{\mu}^{2}(\varpi)$. This yields
(5.6) $\big((P_{\mu}-{\widetilde{P}_{\mu}})f\big|(P_{\mu}-{\widetilde{P}_{\mu}})f\big)_{\varpi}=\big((P_{\mu}-{\widetilde{P}_{\mu}})f^{\perp}\big|(P_{\mu}-{\widetilde{P}_{\mu}})f^{\perp}\big)_{\varpi}$
$=$ $({\widetilde{P}_{\mu}}f^{\perp}|{\widetilde{P}_{\mu}}f^{\perp})_{\varpi}=({\widetilde{P}_{\mu}}f^{\perp}-f^{\perp}|{\widetilde{P}_{\mu}}f^{\perp})_{\varpi},$
where we at the end used $(f^{\perp}|{\widetilde{P}_{\mu}}f^{\perp})_{\varpi}=0$, which follows from ${\widetilde{P}_{\mu}}f^{\perp}\in A_{\mu}^{2}(\varpi)$ and $f^{\perp}\in A_{\mu}^{2}(\varpi)^{\perp}$. We will soon show that
(5.7) $\big|({\widetilde{P}_{\mu}}f^{\perp}-f^{\perp}|{\widetilde{P}_{\mu}}f^{\perp})_{\varpi}\big|\leq C|\eta-\mu|^{\beta}\|f\|_{\varpi}^{2}.$
Combining (5.6)–(5.7) yields $\|P_{\mu}-{\widetilde{P}_{\mu}}\|_{L^{2}(\varpi)\to L^{2}(\varpi)}$ $\leq C|\eta-\mu|^{\beta/2}$, which proves the result, since we obtain from (5.4), (5.5)
$\|P_{\eta}-P_{\mu}\|_{L^{2}(\varpi)\to L^{2}(\varpi)}\leq\|P_{\eta}-{\widetilde{P}_{\mu}}\|_{L^{2}(\varpi)\to L^{2}(\varpi)}+\|{\widetilde{P}_{\mu}}-P_{\mu}\|_{L^{2}(\varpi)\to L^{2}(\varpi)}$
$\leq$ $\|(I_{\varpi}-J_{\mu,\eta}^{-1})P_{\eta}\|_{L^{2}(\varpi)\to L^{2}(\varpi)}+\|J_{\mu,\eta}^{-1}P_{\eta}(I_{\varpi}-J_{\mu,\eta})\|_{L^{2}(\varpi)\to L^{2}(\varpi)}+C|\eta-\mu|^{\beta/2}$
$\leq$ $C^{\prime}|\eta-\mu|^{\beta/2}.$
To see (5.7), (5.5) implies
$\|P_{\eta}J_{\mu,\eta}f-{\widetilde{P}_{\mu}}f\|_{\varpi}\leq C|\eta-\mu|^{\beta}\|f\|_{\varpi}$
for all $f\in L^{2}(\varpi)$. This and (5.4) again yield for $f\in L^{2}(\varpi)$
$\big|\big(f-{\widetilde{P}_{\mu}}f\big|{\widetilde{P}_{\mu}}f\big)_{\varpi}\big|$
$\leq$ $\Big|\big(f-J_{\mu,\eta}f+P_{\eta}J_{\mu,\eta}f-{\widetilde{P}_{\mu}}f\big|{\widetilde{P}_{\mu}}f\big)_{\varpi}\Big|+\Big|\big(J_{\mu,\eta}f-P_{\eta}J_{\mu,\eta}f\big|{\widetilde{P}_{\mu}}f-P_{\eta}J_{\mu,\eta}f\big)_{\varpi}\Big|$
$\leq$ $\Big|\big(f-J_{\mu,\eta}f\big|{\widetilde{P}_{\mu}}f\big)_{\varpi}\Big|+\Big|\big(P_{\eta}J_{\mu,\eta}f-{\widetilde{P}_{\mu}}f\big|{\widetilde{P}_{\mu}}f\big)_{\varpi}\Big|+C|\eta-\mu|^{\beta}\|f\|_{\varpi}^{2}$
$\leq$ $C^{\prime}|\eta-\mu|^{\beta}\|f\|_{\varpi}^{2}.\ \ \Box$
Continuing the proof of Theorem 5.1, the proof of Lemma 4.2 of [11] shows that the set $\Sigma:=\bigcup_{\eta\in{\bf Q}}\sigma(T_{a,\eta})$ is closed. Namely, if the claim were not true, one could find $\lambda\in{\rm cl}\,(\Sigma)\setminus\Sigma$ and sequences $(\eta_{k})_{k=1}^{\infty}\subset{\bf Q}$ and $(\lambda_{k})_{k=1}^{\infty}$ such that $\lambda_{k}\in\sigma(T_{a,\eta_{k}})$ and $\lambda_{k}\to\lambda$ as $k\to\infty$. By passing to a subsequence, if necessary, one may assume that $\eta_{k}\to\eta$ for some $\eta\in{\bf Q}$ as $k\to\infty$. Moreover, since $\sigma(T_{a,\eta})$ is closed, there is a number $0<\delta<1$ such that ${\rm dist}\,(\lambda,\sigma(T_{a,\eta}))\geq\delta\ (*).$ Then, one observes that, due to Lemma 5.3,
(5.8) $\|T_{a,\eta_{k},\eta}-T_{a,\eta_{k}}\|_{A^{2}(\varpi)\to A^{2}(\varpi)}\leq C|\eta-\eta_{k}|^{1/2}.$
Moreover, the spectra of $T_{a,\eta}$ in $A_{\eta}^{2}(\varpi)$ and $T_{a,\eta_{k},\eta}$ in $A_{\eta_{k}}^{2}(\varpi)$ coincide (cf. Lemma $3.$ of [11]), which implies that, for a large enough $k$, $\lambda_{k}$ belongs to the resolvent set of $T_{a,\eta_{k},\eta}$, see ()(*). This eventually leads to a contradiction with $\lambda_{k}\in\sigma(T_{a,\eta_{k}})$ by using some quite standard resolvent estimates and the closeness of the two operators given by formula (5.8).
The closedness of $\Sigma$ implies that $\sigma(T_{a})\subset\Sigma$. To see this, we fix a number $\lambda\in{\mathbb{C}}$ with $\lambda\notin\Sigma$. Then, for every $\eta\in{\bf Q}$, there exists a bounded inverse $R_{\eta,\lambda}:A_{\eta}^{2}(\varpi)\to A_{\eta}^{2}(\varpi)$ of the operator $T_{a,\eta}-\lambda I_{\varpi}$. Since the operator norm $\|R_{\eta,\lambda}\|_{A_{\eta}^{2}(\varpi)\to A_{\eta}^{2}(\varpi)}$ depends continuously on $\eta$ (see the proof of Corollary 4.3 in [11]), it has a uniform upper bound for all $\eta\in{\bf Q}$. We conclude that the operator ${\mathcal{R}}_{\lambda}:f(\cdot,\eta)\mapsto R_{\eta,\lambda}f(\cdot,\eta)$ is bounded in $L^{2}({\bf Q};A^{2}_{\eta}(\varpi))$, and it is the inverse of the operator ${\mathcal{T}}_{a}-\lambda{\mathcal{I}}$ in the space $L^{2}({\bf Q};A^{2}_{\eta}(\varpi))$, where $\mathcal{I}$} is the indentity operator on $L^{2}({\bf Q};A^{2}_{\eta}(\varpi))$. Hence, ${\sf F}^{-1}{\mathcal{R}}_{\lambda}{\sf F}$ is a bounded inverse of $T_{a}-\lambda I_{\Omega}={\sf F}^{-1}({\mathcal{T}}_{a}-\lambda{\mathcal{I}}){\sf F}$, and $\lambda$ thus belongs to the resolvent set of $T_{a}$. We refer to [11] for the details of these arguments.
To complete the proof of Theorem 5.1, it is thus sufficient to show that $\Sigma\subset\sigma_{\rm ess}(T_{a})$, i.e. every $\lambda$, which belongs to $\sigma(T_{a,\mu})$ for some $\mu\in{\bf Q}$, is a point in the essential spectrum of $T_{a}$. To this end, we will use the Weyl criterion as presented e.g. in Lemma $1.$ of [11], and thus we fix $\mu\in{\bf Q}$, $\lambda\in\sigma(T_{a,\mu})$, an arbitrary $\varepsilon>0$ and a near eigenfunction $g\in A_{\mu}^{2}(\varpi)$ with $\|g\|_{\varpi}=1$ such that
(5.9) $\|T_{a,\mu}g-\lambda g\|_{\varpi}\leq\varepsilon.$
Since the operator $T_{a,\mu}$ is bounded and $A_{\mu,{\rm ext}}^{2}(\varpi)$ is dense in $A_{\mu}^{2}(\varpi)$, it can be assumed that $g\in A_{\eta,{\rm ext}}^{2}(\varpi)$.
Given $n\in{\mathbb{N}}$, we define $G_{n}={\sf F}^{-1}({\mathcal{X}}_{\mu,n}g),$ where $({\mathcal{X}}_{\mu,n}g)(z,\eta):=g\otimes{\mathcal{X}}_{\mu,n}(z,\eta)=g(z){\mathcal{X}}_{\mu,n}(\eta)\in L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$ with
(5.10) ${\mathcal{X}}_{\mu,n}(\eta)=\left\{\begin{array}[]{ll}n,&\eta\in B(\mu,1/n)\\ 0,&\mbox{for other}\ \eta\in{\bf Q},\end{array}\right$.
where $B(\mu,1/n)=\{\eta\in{\bf Q}\,:\,|\eta-\mu|<1/n\}$. We have $1/2\leq\int_{\bf Q}{\mathcal{X}}_{\mu,n}(\eta)^{2}d\eta\leq 4$ for all $n$, hence
$1/2\leq\|{\mathcal{X}}_{\mu,n}g\|_{L^{2}({\bf Q};L^{2}(\varpi))}^{2}=\int\limits_{\bf Q}{\mathcal{X}}_{\mu,n}(\eta)^{2}\|g\|_{\varpi}^{2}d\eta\leq 4.$
Taking the Floquet inverse transform yields $G_{n}={\sf F}^{-1}({\mathcal{X}}_{\mu,n}g)\in A^{2}(\Omega)$ and
$1/2\leq\|G_{n}\|_{\Omega}\leq 4.$
Using Lemma 2.4 we choose for every $n$ the number $\rho(n)>0$ such that $\rho(n+1)<\rho(n)$ and
(5.11) $\|G_{n}-\psi^{n}G_{n}\|_{\Omega}\leq\frac{1}{n+2}\ \ \ \mbox{with}\ \ \ \psi^{n}:=\psi^{(\rho(n))}.$
The Weyl singular sequence for the number $\lambda$ is defined by the translated functions $h_{n}=f_{n}\circ\mathsf{t}_{n}\|f_{n}\|_{\Omega}^{-1}$, where $f_{n}=\psi^{n}G_{n}\in A^{2}(\Omega)$ and $\mathsf{t}_{n}(z)=z-m(n)$, and the numbers $m(n)$ are chosen so as to form a fast enough increasing sequence of positive integers. Note that the above definitions imply $1/C^{\prime}\leq\|f_{n}\|_{\Omega}\leq C^{\prime}$ for some constant $C^{\prime}>0$, for all $n$, hence, it is plain that condition (1.4) of Lemma 1.1 in [11] holds, once we verify the estimate
(5.12) $\|T_{a}f_{n}-\lambda f_{n}\|_{\Omega}\leq C\varepsilon$
for all large enough $n\in{\mathbb{N}}$ (here, the constant $C$ may depend on $\lambda$). To see this, we assume that $n\geq\varepsilon^{-2/\beta}$ in the following. Since $\|T_{a}(f_{n}-G_{n})\|_{\Omega}<C\varepsilon$ by (5.11) and the boundedness of the operator $T_{a}$ in $L^{2}(\Omega)$, it suffices to show that $\|T_{a}G_{n}-\lambda G_{n}\|_{\Omega}<C\varepsilon.$
We write
$T_{a}G_{n}$ $=$ ${\sf F}^{-1}{\mathcal{P}}{\mathcal{M}}_{a}({\mathcal{X}}_{\mu,n}g)={\sf F}^{-1}\big({\mathcal{X}}_{\mu,n}(\eta)P_{\eta}(ag)(z)\big)$
$=$ ${\sf F}^{-1}\Big({\mathcal{X}}_{\mu,n}(\eta)\big(P_{\eta}(ag)(z)-P_{\mu}(ag)(z)\big)\Big)$
$\ \ +$ ${\sf F}^{-1}\big({\mathcal{X}}_{\mu,n}(\eta)P_{\mu}(ag)(z)-\lambda{\mathcal{X}}_{\mu,n}(\eta)g(z)\big)+{\sf F}^{-1}\big(\lambda{\mathcal{X}}_{\mu,n}g\big)$
$=:$ $\Psi_{1}+\Psi_{2}+\lambda G_{n}.$
Lemma 5.3 yields the bound $\|P_{\eta}-P_{\mu}\|_{A^{2}(\varpi)\to A^{2}(\varpi)}\leq C|\eta-\mu|^{\beta/2}$. We get by (5.10)
(5.13) $\big\|{\mathcal{X}}_{\mu,n}\big(P_{\eta}(ag)-P_{\mu}(ag)\big)\big\|_{L^{2}({\bf Q};L^{2}(\varpi))}^{2}$
$\leq$ $C_{a}\!\!\!\!\!\int\limits_{B(\mu,1/n)}\!\!\!\!\!\!n^{2}|\eta-\mu|^{\beta}\|g\|_{\varpi}^{2}d\eta\leq\frac{C^{\prime}_{a}}{n^{\beta}}\leq C^{\prime}_{a}\varepsilon^{2}.$
The near eigenfunction property (5.9) yields a similar estimate for $\Psi_{2}$:
$\big\|{\mathcal{X}}_{\mu,n}\big(P_{\mu}(ag)-\lambda g\big)\big\|_{L^{2}({\bf Q};L^{2}(\varpi))}^{2}\leq C\!\!\!\!\!\!\int\limits_{B(\mu,1/n)}\!\!\!\!\!\!n^{2}\varepsilon^{2}d\eta\leq C^{\prime}\varepsilon^{2}.$
From this and (5.13) we obtain $\|\Psi_{1}\|_{\Omega}+\|\Psi_{2}\|_{\Omega}\leq C\varepsilon$, and this shows that the functions $h_{n}$ satisfy (5.12) and thus (1.4) of Lemma 1.1 in [11].
There remains to show that the sequence $(h_{n})_{n=1}^{\infty}$ does not have subsequences converging in $L^{2}(\Omega)$. This is intuitively quite clear, since the functions $f_{n}$ have some localization in $\Omega$ due to the factors $\psi^{n}$. Choosing a rapidly enough growing sequence $(m(n))_{n=1}^{\infty}$ for the definition of the translations $\mathsf{t}_{n}$, see above, one can arrange that $\|h_{n}-h_{m}\|_{\omega}$ has a positive lower bound independent of $n,m$, since $h_{n}$ is essentially the same as $f_{n}\circ\mathsf{t}_{n}$. The details of the proof can be completed by replacing $\varphi_{n}$ by $\psi^{n}$ and by other obvious changes at the end of Section 4 of [11]. This completes the proof of Theorem 5.1.

### v2

@@
 The research of the first named author was partially supported by the Academy of Finland grant no. 359563
+Addresses:
+1 Department of Mathematics and Statistics, University of Helsinki, P.O. Box 68, 00014 Helsinki, Finland
+email jari.taskinen@helsinki.fi
+2 Department of Mathematics and Statistics, University of Helsinki, P.O. Box 68, 00014 Helsinki, Finland
+email zhan.zhang@helsinki.fi
 1. Introduction.
@@
 We will apply these considerations to study the spectra of Toeplitz operators with doubly periodic symbols.
-Given a domain $\Omega$ in the complex plane $\mathbb{C}$}, we denote by $L^{2}(\Omega)$ the usual Lebesgue-Hilbert space with respect to the (real) area measure dAdA and by $A^{2}(\Omega)$ the corresponding Bergman space, which is the closed subspace consisting of analytic functions.
+Given a domain $\Omega$ in the complex plane $\mathbb{C}$}, we denote by $L^{2}(\Omega)$ the usual Lebesgue-Hilbert space with respect to the (real) area measure $dA$ and by $A^{2}(\Omega)$ the corresponding Bergman space, which is the closed subspace consisting of analytic functions.
 The Floquet transform is defined for $f\in A^{2}(\Omega)$ by
@@
 However, we will pose one more assumption, related to the theory of elliptic functions, see $({\mathcal{A}})$ in the next section.
-Here and later, cl(A)(A) denotes the closure of the set AA.
-Definition $1.$
+Here and later, cl(A)(A) denotes the closure of the set $A$.
+Definition 1.1.
 We denote by $\varpi$ the periodic cell, which is a domain in $\mathbb{C}$} such that
@@
 We emphasize that contrary to [10], we do not need to add assumptions on the smoothness of the boundary $\partial\Omega$.
-Anyway, the periodic domain $\Omega$ is always infinitely connected, due to assumption $(iii)$ of Definition $1.$
+Anyway, the periodic domain $\Omega$ is always infinitely connected, due to assumption $(iii)$ of Definition 1.1.
 The following notation will be used throughout the paper.
@@
 If $x$ is a point in $\mathbb{C}$} or ${\mathbb{R}}^{n}$ and $r>0$, then $B(x,r)$ denotes the Euclidean ball with center $x$ and radius $r>0$.
-Moreover, cl(A)(A) denotes the closure of a set AA.
+Moreover, cl(A)(A) denotes the closure of a set $A$.
 Given a domain $D\subset{\mathbb{C}}$, we denote by $\|f\|_{D}$ and $(\cdot|\cdot)_{D}$ the norm and inner product of $L^{2}(D)$, which is the $L^{2}$ space with respect to the (real) area measure.
 In general, the norm of a Banach space $X$ is denoted by $\|\cdot\|_{X}$.
-Given an interval $I\subset{\mathbb{R}}$ and a Banach space $X$ we denote by $L^{2}(I;X)$ the space of vector valued, Bochner-$L^{2}$-integrable functions on II with values in $X$, endowed with the norm
+Given an interval $I\subset{\mathbb{R}}$ and a Banach space $X$ we denote by $L^{2}(I;X)$ the space of vector valued, Bochner-$L^{2}$-integrable functions on $I$ with values in $X$, endowed with the norm
 $\|f\|_{L^{2}(I;X)}=\Big(\int\limits_{I}\|f(t)\|_{X}dt\Big)^{1/2}\,.$
@@
 209) and simple functions of $L^{2}\big({\bf Q};A^{2}(\varpi)\big)$ are not in general contained in these subspaces.
-Lemma $3.$
+Lemma 3.3.
 The space ${\mathcal{H}}_{\eta}$ is a dense subspace of $L^{2}\big({\bf Q};A_{\eta}^{2}(\varpi)\big)$.
@@
 We first observe that by Proposition 3.2.(ii)(ii), $\sf F$} maps the dense subspace $X$ of $A^{2}(\Omega)$ into ${\mathcal{H}}_{\eta}$.
-In view of the unitarity of $\sf F$}, this implies that ${\sf F}(A^{2}(\Omega))$ is contained in the closure of ${\mathcal{H}}_{\eta}$ in $L^{2}({\bf Q};L^{2}(\varpi))$, which equals $L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$, by Lemma $3.$
+In view of the unitarity of $\sf F$}, this implies that ${\sf F}(A^{2}(\Omega))$ is contained in the closure of ${\mathcal{H}}_{\eta}$ in $L^{2}({\bf Q};L^{2}(\varpi))$, which equals $L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$, by Lemma 3.3.
 We are thus left with the construction of an inverse image under $\sf F$} for all $g\in L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$.
@@
 Moreover, the spectra of $T_{a,\eta}$ in $A_{\eta}^{2}(\varpi)$ and $T_{a,\eta_{k},\eta}$ in $A_{\eta_{k}}^{2}(\varpi)$ coincide (cf.
-Lemma $3.$ of [11]), which implies that, for a large enough $k$, $\lambda_{k}$ belongs to the resolvent set of $T_{a,\eta_{k},\eta}$, see ()(*).
+Lemma 3.3.
+of [11]), which implies that, for a large enough $k$, $\lambda_{k}$ belongs to the resolvent set of $T_{a,\eta_{k},\eta}$, see ()(*).
 This eventually leads to a contradiction with $\lambda_{k}\in\sigma(T_{a,\eta_{k}})$ by using some quite standard resolvent estimates and the closeness of the two operators given by formula (5.8).
@@
 To this end, we will use the Weyl criterion as presented e.g.
-in Lemma $1.$ of [11], and thus we fix $\mu\in{\bf Q}$, $\lambda\in\sigma(T_{a,\mu})$, an arbitrary $\varepsilon>0$ and a near eigenfunction $g\in A_{\mu}^{2}(\varpi)$ with $\|g\|_{\varpi}=1$ such that
+in Lemma 1.1.
+of [11], and thus we fix $\mu\in{\bf Q}$, $\lambda\in\sigma(T_{a,\mu})$, an arbitrary $\varepsilon>0$ and a near eigenfunction $g\in A_{\mu}^{2}(\varpi)$ with $\|g\|_{\varpi}=1$ such that
 (5.9) $\|T_{a,\mu}g-\lambda g\|_{\varpi}\leq\varepsilon.$